A dilute aqueous CaCl2 solution at 25 degrees celsius has an osmotic pressure pi = 10bar. What is the vapor pressure(in bar units) above the solution? The density of water at 25 degrees Celsius is water (25 degrees Celsius) = 0.997gmL.
Answer #1

First, the osmotic pressure in atmospheres is calculated:

P = 10 bar * (1 atm / 1,01325 bar) = 9.87 atm

The concentration of ions in solution is calculated:

C = P / R * T = 9.87 atm / 0.082 atm * L / mol * K * 298 K = 0.40 M

The molar fraction of ions in solution is calculated:

X i = 0.40 mol / L * (1 L / 1000 mL) * (1 mL / 0.997 g) * (18 g / 1 mol) = 0.007

The vapor pressure of the solution is calculated, for a water vapor pressure of 23.776 mmHg at 25 ° C:

Pd = Pvap * (1 - X i) = 23.776 mmHg * (1 - 0.007) = 23.610 mmHg

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.